Need help as soon as possible please!!

Crane, Inc. buys 1 CD 19 Determine the cost per CD for each inal answers to 2 declmal places, e.g. 5.8.) Cost per CD based on the Point Potential Policy set by your instructor. Question Attempts: O of 6 used SAVE POR LATER 6 8 9 O.
Answer #1

Calculate cost per CD under relative sales value method

Group Relative sale value Total Cost allocated Cost per CD
1 200*6 = 1200 9500*1200/15400 = 740.26 740.26/200 = 3.70
2 800*13 = 10400 9500*10400/15400 = 6415.58 6415.58/800 = 8.02
3 200*19 = 3800 9500*3800/15400 = 2344.16 2344.16/200 = 11.72
Total 15400 9500